# On the polemic assessment of what Bell did

### v2

@@ III.1.2 Quantum nonlocality
In [6], Bell gave a qualitative description of why a locally causal explanation is absent and concluded that in quantum mechanics “We simply do not have (2)”, where (2) in his paper is LC.
It is relevant that by locally causal explanation he meant any explanation, not only “real” explanations in the sense of preexisting entities. His argument is similar to the one given by Einstein in 1927.666Einstein’s argument is explained by Laudisa [22] and also by Harrigan and Spekkens [23]. It is a genuine nonlocality argument that does not contain speculations about infamous “elements of physical reality.” Both Einstein and Bell were concerned with the existence of “actions” or “influences” at a distance, not with the need for preexisting entities or with denying quantum superposition, which are subsidiary, irrelevant issues regarding their actual arguments. [^7] Those issues were obfuscated in the EPR paper by the “elements of physical reality” concept, which neither Einstein nor Bell used in their arguments [24].
@@
Given that (5) is not widely known as a proof of quantum nonlocality, appendix A further explains the local causality concept, and appendix A.2 clarifies a common misunderstanding claiming to refute (5) via a counterexample.
Also, some reject (4) as incorrect because it would somehow conflict with quantum “nonrealism.” However, according to Bell, “…nothing is said about the locality, or even localizability, of the variable $\lambda$. These variables could well include, for example, quantum mechanical state vectors, which have no particular localization in ordinary space-time…” [18]. Thus, in Bell’s view, and lining up with Reichenbach’s Common Cause Principle, it is unjustified to restrict $\lambda$ to be an “element of physical reality” in the EPR sense (cf. appendix B).
@@
III.1.3 The gist of the endemic controversy
@@
For some strange reason, those who argue quantum mechanics is nonlocal dismiss (7) as proof despite, in some cases, being well aware of its existence. They prefer to turn to the Bell inequality as their sole argument, conflating the issues of quantum nonlocality with the possibility of a local completion [15], something that, as we have seen, Bell was careful to distinguish.
On the other hand, those claiming that quantum mechanics is local rely on point 1, conveniently blocking the “usual” quantum nonlocality proof.
III.2 Statistical Independence
@@
III.3 The EPR Paper
It is remarkable that in his 1975 article “The theory of local beables” [6], Bell did not reference the EPR paper. As noted in section II.1, in 1964, Bell conceived his article as a continuation of the EPR paper, assuming quantum nonlocality to investigate the possibility of a local completion.
@@ III.3 The EPR Paper
Perhaps motivated by the orthodox criticism against the EPR reasoning, unlike in his 1964 approach, the above (12) quantum nonlocality proof was not implicitly assumed in his 1975 paper, which is why he introduced the concept of local causality to provide an independent and explicit proof. The proof was directly applicable to nondeterministic theories such as quantum mechanics, eliminating the realism tenet that, to Einstein’s dislike, somehow obscured the EPR argument [24].
@@ IV Conclusions
Unlike the conventional interpretation, a careful analysis of Bell’s papers reveals that he considered his theorem to be about the impossibility of a local completion of quantum mechanics, rather than a direct proof of quantum nonlocality which, according to him, demanded an independent proof.
Logical rigor suggests following Bell’s approach; otherwise, an argumentative loophole may arise: if we do not first prove that orthodox quantum mechanics is nonlocal, then the fact that a different, more complete modified theory is nonlocal does not necessarily imply the nonlocality of the first one.
@@
A rational alternative for arguing quantum locality could be to reject local causality as a valid criterion of locality and postulate that quantum mechanics is local. Given its no-signaling property, this is logically and empirically admissible, but a seldom, perhaps never, explicitly admitted path. Instead, flawed arguments are presented alongside obscure speculations about realism, thereby exacerbating the endemic controversy.
Apéndice A Local Causality
Here we present a more detailed analysis of the local causality concept and give an example of a common misunderstanding. We refer the reader to the excellent exposition of local causality by Norsen [15].
A.1 The LC concept
Bell’s definition of LC is a formalization of the idea that, according to relativity theory, interactions can happen only at a finite speed. It means that causes cannot have an instantaneous effect on distant events. He formulated LC so that it can be applied to not deterministic theories like quantum mechanics. It is a locality argument that avoids a purportedly classical EPR-like reasoning. A concept directly applicable to orthodox quantum mechanics without distorting its nature.
@@ A.1 The LC concept
The common cause $\lambda$ is usually called “hidden variables”; however, it is somewhat misleading to believe the $\lambda$ variables are necessarily unknown parameters. The only condition they need to comply with is lying at the intersection of the backward light cones of the measuring events to constitute a local explanation of the correlations. Appendix B further expands on the meaning of the common causes.
A.2 An alleged counterexample disproving LC
@@
[…]
@@
[…]
@@
Apéndice B Common Causes meaning
@@
Sustaining that $\lambda$, understood as a common cause, is by necessity an element foreign to quantum mechanics amounts to forbidding the application of the local causality concept to quantum mechanics, that according to Bell’s reasoning would be unjustified. The general nature of the $\lambda$ variables and that they may well include the quantum state was explicitly mentioned by Bell[18]:
@@
Apéndice C Proof of formula (5)
We can test if the quantum state can be a local common cause explaining the perfect anti-correlations by setting $\lambda=\mid\psi\rangle$ in (19) with $\mid\psi\rangle$ given by (3), where $\mid+\rangle$ and $\mid$-$\rangle$ denote the spin eigenstates in the $z$-direction. We assume that motion takes place in the $y$ direction with setting angles $a$ and $b$ lying in the xzx-z plane measured with respect to the $z$ axis. If $\mid a,+\rangle$ and $\mid a,$-$\rangle$ are the spin eigenstates in the $a$ direction.

### v3

@@ 3.1.2 Quantum nonlocality
In [Bell, Shimony, Horne, and Clauser 1985], Bell gave a qualitative description of why a locally causal explanation is absent and concluded that in quantum mechanics “We simply do not have (2)”, where (2) in his paper is LC.
It is relevant to realize that by locally causal explanation he meant any explanation, not only “real” explanations in the sense of preexisting entities. His argument is similar to the one given by Einstein in 1927.777Einstein’s argument is explained by Laudisa [Laudisa 2019] and also by Harrigan and Spekkens [Harrigan and Spekkens 2010]. It is a genuine nonlocality argument that does not contain speculations about the infamous “elements of physical reality.” Both Einstein and Bell were concerned with the existence of “actions” or “influences” at a distance. They were not concerned with the need for preexisting entities, determinism, or the denial of quantum superposition that are commonly associated with the notion of realism. Those were subsidiary and irrelevant issues regarding their true concerns about nonlocality. [^8] Those issues were obfuscated in the EPR paper by the “elements of physical reality” concept, which neither Einstein nor Bell used in their arguments [Howard 1985].
@@
Given that (5) is not widely known as a proof of quantum nonlocality, appendix 5.1 further explains the local causality concept, and appendix 5.2.1 clarifies a common misunderstanding claiming to refute (5) via a counterexample.
Also, some reject (4) as incorrect because it would somehow conflict with quantum “nonrealism.” However, according to Bell, “…nothing is said about the locality, or even localizability, of the variable $\lambda$. These variables could well include, for example, quantum mechanical state vectors, which have no particular localization in ordinary space-time…” [Bell 1981]. Thus, in Bell’s view, and lining up with Reichenbach’s Common Cause Principle, it is unjustified to restrict $\lambda$ to be an “element of physical reality” in the EPR sense (cf. appendix 5.3).
@@
3.1.3 The gist of the endemic controversy
@@
For some strange reason, those who argue quantum mechanics is nonlocal dismiss (7) as proof despite, in some cases, being well aware of its existence. They prefer to turn to the Bell inequality as their sole argument, conflating the issues of quantum nonlocality with the possibility of a local completion [Norsen 2011], something that, as we have seen, Bell was careful to distinguish.
On the other hand, those claiming that quantum mechanics is local rely on one of the following strategies:
  • recursing to point 1) above, conveniently blocking the “usual” quantum nonlocality proof.
  • invoking realism or the “elements of physical reality” idea that is universally attached to the EPR reasoning.
The first point above is arguably well justified for rejecting quantum nonlocality. However, the second one, although very popular, lacks rational support because Bell’s local causality was conceived independently of any notion of determinism or preexisting values and cannot be consistently applied to this case.
3.2 Statistical Independence
@@
3.3 The EPR Paper
Although in 1964, Bell conceived his article as a continuation of the EPR paper (assuming quantum nonlocality to investigate the possibility of a local completion), it should not be surprising that in his 1975 paper “The theory of local beables” [Bell, Shimony, Horne, and Clauser 1985], Bell did not reference the EPR manuscript. This time, his intention was to provide an argument independently of EPR and determinism.
@@ 3.3 The EPR Paper
Perhaps motivated by the orthodox criticism against the EPR reasoning, unlike in his 1964 approach, the above (12) quantum nonlocality proof was not implicitly assumed in his 1975 paper, which is why he introduced the concept of local causality to provide an independent proof. The proof was directly applicable to non-deterministic theories such as quantum mechanics, eliminating the realism tenet that, to Einstein’s dislike, somehow obscured the EPR argument [Howard 1985].
@@ 4 Conclusions
Unlike the conventional interpretation, a careful analysis of Bell’s papers reveals that he considered his theorem to be, as he once put it, “… a theorem of mine about hidden variables” [Bell 1975], rather than a theorem about quantum nonlocality which, according to him, demanded an independent proof.
Logical rigor suggests following Bell’s approach; otherwise, an argumentative loophole may arise: if we do not first prove that orthodox quantum mechanics is nonlocal, then the fact that a different modified theory is nonlocal does not necessarily imply the nonlocality of the first one.
@@
A rational alternative for arguing quantum locality could be to reject local causality as a valid criterion of locality and postulate that quantum mechanics is local. Given its no-signaling property, this is logically and empirically admissible, but a seldom, perhaps never, explicitly admitted path. Instead, obscure and unconvincing metaphysical speculations about realism are rehearsed, which are logically disconnected from the nonlocality problem, thereby exacerbating the existing controversy [Norsen 2007, Gisin 2012].
A probably more drastic alternative is to embrace superdeterminism and consider quantum mechanics as an emergent theory [’t Hooft 2021].
5 APPENDIX
5.1 Local Causality
Here we present a more detailed analysis of the local causality concept and give an example of a common misunderstanding. We refer the reader to the excellent exposition of local causality by Norsen [Norsen 2011].
5.2 The LC concept
Bell’s definition of LC is a formalization of the idea that, according to relativity theory, interactions can happen only at a finite speed. It means that causes cannot have an instantaneous effect on distant events. He formulated LC so that it can be applied to nondeterministic theories like quantum mechanics. It is a locality argument that avoids a purportedly classical EPR-like reasoning. A concept directly applicable to orthodox quantum mechanics without distorting its nature.
@@ 5.2 The LC concept
The common cause $\lambda$ is usually called “hidden variables”; however, it is somewhat misleading to believe the $\lambda$ variables are necessarily unknown parameters. The only condition they need to comply with is lying at the intersection of the backward light cones of the measuring events to constitute a local explanation of the correlations. Appendix 5.3 further expands on the meaning of the common causes.
5.2.1 An alleged counterexample disproving LC
@@
$\left.\begin{array}[]{ccccc}P(R,B\mid L1,L2)&=&P(B,R\mid L1,l2)&=&1/2\\ P(R,R\mid L1,L2)&=&P(B,B\mid L1,l2)&=&0\\ P(R\mid L1)&=&P(B\mid L1)&=&1/2\\ P(R\mid L2)&=&P(B\mid L2)&=&1/2\end{array}\right\}$ (20)
@@
$\left.\begin{array}[]{ccc}P(R,B\mid L1,L2,\lambda=(R,B))&=&1\\ P(B,R\mid L1,L2,\lambda=(R,B))&=&0\\ P(R\mid L1,\lambda=(R,B))&=&1\\ P(B\mid L1,\lambda=(R,B))&=&0\\ P(R\mid L2,\lambda=(R,B))&=&0\\ P(B\mid L2,\lambda=(R,B))&=&1\end{array}\right\}$ (22)
@@
5.3 Common Causes meaning
@@
Sustaining that $\lambda$, understood as a common cause, is by necessity an element foreign to quantum mechanics amounts to forbidding the application of the local causality concept to quantum mechanics, that according to Bell’s reasoning would be unjustified. The general nature of the $\lambda$ variables and that they may well include the quantum state was explicitly mentioned by Bell[Bell 1981]:
@@
5.4 Proof of formula (5)
We can test if the quantum state can be a local common cause explaining the perfect anti-correlations by setting $\lambda=\mid\psi\rangle$ in (19) with $\mid\psi\rangle$ given by (3), where $\mid+\rangle$ and $\mid$-$\rangle$ denote the spin eigenstates in the $z$-direction. We assume that motion takes place in the $y$ direction with setting angles $a$ and $b$ lying in the $xz$ plane measured with respect to the $z$ axis. If $\mid a,+\rangle$ and $\mid a,$-$\rangle$ are the spin eigenstates in the $a$ direction.